# Trees of Graphs as Boundaries of Hyperbolic Groups

### v1

@@ 4.1. Trees of graphs as inverse limits of graphs
induced by the map $\varphi_{[t,u]}\circ\varphi^{-1}_{[s,u]}:\varphi_{[s,u]}(\Sigma_{u})\to\varphi_{[t,u]}(\Sigma_{u})$ (under the canonical identification of each of the two divisors with the corresponding peripheral subspace), and that $b_{t,s}=b_{s,t}^{-1}$. The equivalence relation $\sim$ is then determined by all of the maps $b_{s,t}$ as above related to all white vertices $u$ of $F$. Note that the so described quotient space $X_{F}$ may be viewed as obtained by appropriate iterated connected sum of the family $\Gamma_{t}:t\in V^{b}_{F}$ of graphs (see the paragraph right before Lemma 4.8 for a precise description of the concept of connected sum for graphs). Note also that $X_{F}$ is in a natural way homeomomorphic to the space $K^{*}_{F}$, via homeomorphism induced by the homeomorphisms $h_{t}^{-1}:|\Gamma_{t}|\to K^{*}_{t}$, for $t\in V^{b}_{F}$.
In order to make our alternative description of the inverse system ${\mathcal{S}}_{\Theta}$ complete, we need to describe the bonding maps between the spaces $X_{F}$ which, under the above identifications with the spaces $K_{F}^{*}$, correspond to the bonding maps $f_{FF^{\prime}}$. To do this, consider any two finite b-subtrees $F\subset F^{\prime}$ of $T$, and define the map $\pi_{FF^{\prime}}:X_{F^{\prime}}\to X_{F}$ to be a natural map from the larger connected sum to the smaller one, given as follows. For $t\in V^{b}_{F}$, the restriction of $\pi_{FF^{\prime}}$ to $|\Gamma_{t}|^{\#}(J_{t}^{F^{\prime}})$ (viewed as a subset in $X_{F^{\prime}}$) coincides with the blow-down map $\rho^{|\Gamma_{t}|}_{J_{t}^{F^{\prime}}J_{t}^{F}}$ (whose image $|\Gamma_{t}|^{\#}(J_{t}^{F})$ is viewed as a subset in $X_{F}$). For $t\in V^{b}_{F^{\prime}}\setminus V^{b}_{F}$, the subset $|\Gamma_{t}|^{\#}(J_{t}^{F^{\prime}})\subset X_{F^{\prime}}$ is mapped by $\pi_{FF^{\prime}}$ to the point $h_{s}([\varphi_{[}u,s](\Sigma_{u})])\in|\Gamma_{s}|^{\#}(J_{s}^{F})\subset X_{F}$, where $s\in V^{b}_{F}$ is the closest to $t$ (in the polygonal metric in $T$) vertex of the subtree $F$, and where $u\in V^{w}_{F^{\prime}}$ is the vertex adjacent to $s$ on the shortest path in $T$ from $t$ to $s$.
@@
Expression of a tree of graphs $\lim\Theta$ as the inverse limit of graphs $X_{F}$, as above, turns out to be very useful for establishing various topological properties of trees of graphs. Next corollary records most obvious such properties (which follow directly from the corresponding well known properties of inverse limits, e.g. the estimate of the topological dimension of the limit by supremum of the topological dimensions of the terms in the inverse system). Further properties, for a subclass of trees of 2-connected graphs, are discussed in the next subsection.
@@
Any tree of graphs is a compact metrisable topological space of topological dimension $\leq 1$.
@@ 4.2. Trees of 2-connected graphs
A graph is 2-connected if it is finite, connected, nonempty, not reduced to a single vertex, and contains no cutpoint. We say that a tree system of punctured graphs $\Theta=(T,\{\Gamma_{t}^{\circ}\},\{\Sigma_{u}\},\{\varphi_{e}\})$ in which all the graphs $\Gamma_{t}$ are 2-connected is a tree system of punctured 2-connected graphs; likewise, a tree of 2-connected graphs is the limit of a tree system of punctured 2-connected graphs.
@@
Let ${\mathcal{S}}=(\{K_{\alpha}\}_{\alpha},\{f_{\alpha\alpha^{\prime}}\}_{\alpha\leq\alpha^{\prime}})$ be an inverse system of compact metric spaces, and let $C$ be a closed subspace of some $K_{\alpha}$. Then the preimage $f_{\alpha}^{-1}(C)$ of $C$ in $Z$ is disconnected if and only if $f_{\alpha\hat{\alpha}}^{-1}(C)$ is disconnected for some $\hat{\alpha}\geq\alpha$.
@@
Let $X$ be a connected, locally connected, cutpoint-free $T_{1}$ space. Let $x,x^{\prime}\in X$ and let $C$ be a component of $X\setminus\{x,x^{\prime}\}$. The closure $\bar{C}$ of $C$ in $X$ is $C\cup\{x,x^{\prime}\}$.
@@
If $x=x^{\prime}$ then we can directly apply Lemma 4.8 so we assume that $x\neq x^{\prime}$. Let $X_{y}=X\setminus\{y\}$. Then $C$ is a component of $X_{y}\setminus\{x\}$ so, Lemma 4.8, the closure of $C$ in $X_{y}$ is $C\cup\{x\}$. Then $\bar{C}$ is either $C\cup\{x\}$ or $C\cup\{x,x^{\prime}\}$. Swapping the roles of $x$ and $x^{\prime}$ we see that $\bar{C}=C\cup\{x,x^{\prime}\}$. ∎
@@
Let $X_{1}$ and $X_{2}$ be realizations graphs, suppose that $X_{2}$ is 2-connected, and let $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$. Let $\ell\colon P_{x_{1}}\to P_{x_{2}}$ be a bijection between the blow-up divisor $P_{x_{1}}$ of $x_{1}$ in $X_{1}$ and the blow-up divisor $P_{x_{2}}$ of $x_{2}$ in $X_{2}$. Let $A$ be a connected subspace of $X_{1}$. Then the preimage of $A$ under the projection $X_{1}\#_{\ell}X_{2}\to X_{1}$ is connected.
@@
It is clear that $X$ is not empty and not reduced to a single point. So we need only show that $X$ is connected and has no cutpoint. That $X$ is connected follows from Lemma 4.10 by setting $A=X_{1}$.
@@
By Fact 4.5, it suffices to prove that $f_{FF^{\prime}}^{-1}(C)$ is connected for any b-subtree $F^{\prime}\supset F$. We prove this by induction on the size of $F^{\prime}\setminus F$. The base case $F^{\prime}=F$ is immediate. Otherwise, the reduced partial union graph $K_{F^{\prime}}^{\ast}$ is the connected sum of the reduced partial union graph $K_{F^{\prime\prime}}^{\ast}$ and the realization of a 2-connected graph, for some b-subtree $F^{\prime\prime}$ with $F^{\prime}\supsetneq F^{\prime\prime}\supset F$. Then $f_{FF^{\prime}}^{-1}(C)=f_{F^{\prime\prime}F^{\prime}}^{-1}(\hat{C})$ where $\hat{C}=f_{FF^{\prime\prime}}^{-1}(C)$, which is connected by the inductive hypothesis. By Lemma 4.11, the graph $K_{F^{\prime\prime}}^{\ast}$ is also 2-connected so, by Lemma 4.10, the preimage $f_{F^{\prime\prime}F^{\prime}}^{-1}(\hat{C})=f_{FF^{\prime}}^{-1}(C)$ is connected. ∎
@@
Let ${\mathcal{S}}_{\Theta}=(\{K_{F}^{\ast}:F\in{\mathcal{F}}_{T}^{b}\},\{f_{FF^{\prime}}:F\subset F^{\prime}\})$ be the standard inverse system associated to $\Theta$. By Lemma 4.11, any reduced partial union $K_{F}^{\ast}$ is connected. Then, since the inverse limit of connected metric compacta is connected, $Z$ is connected as well.
@@
whose limit $\lim_{\leftarrow}{\mathcal{C}}$ is obviously a subspace of $Z$. Observe that, since $\mathcal{C}$ is the sequence of circles, with bonding maps which are near-homeomorphisms (i.e. can be approximated by homeomorphisms), by the result of M. Brown [3] the limit $\lim_{\leftarrow}{\mathcal{C}}$ is also a circle, which completes the proof. ∎
5. Necessary condition for a hyperbolic group to have a tree of graphs as its Gromov boundary
[…]
@@ 5. Necessary condition for a hyperbolic group to have a tree of graphs as its Gromov boundary
Let $X$ be a connected, locally connected, cutpoint-free, compact Hausdorff space. An infinite subspace $B\subset X$ is a subnecklace if it satisfies the following conditions.
@@
    For any $p,q\in B$, the space $X\setminus\{p,q\}$ has two components.
@@
    For any $p\in B$, the space $X\setminus\{p\}$ has two ends (i.e. $p$ is a local cutpoint of degree 2).
@@
We now turn to the analysis of subnecklaces in connected, locally connected and cutpoit-free trees of graphs. We start with terminological preparations, and with some preparatory technical observations. Let $\Theta=(T,\{\Gamma_{t}^{\circ}\},\{\Sigma_{u}\},\{\varphi_{e}\})$ be a tree system of punctured graphs, let ${\mathcal{S}}_{\Theta}=(\{K_{F}^{\ast}:F\in{\mathcal{F}}_{T}^{b}\},\{f_{FF^{\prime}}:F\subset F^{\prime}\})$ be the standard inverse system associated to $\Theta$ and assume that the limit $Z=\lim\Theta=\lim_{\leftarrow}\mathcal{S}_{\Theta}$ is connected and locally connected. As described in Section 4.1, for a finite b-subtree $F$ of $T$, the reduced partial union $K_{F}^{\ast}=K_{F}/\Omega_{F}$ of $\Theta$ relative to $F$ is a graph. Let $D_{F}$ be the image of the peripheral subspaces $\Omega_{F}$ in $K_{F}^{\ast}$ and note that $D_{F}$ is dense, countable and contains the vertex set of $K_{F}^{\ast}$. Note that any $p\in K_{F}^{\ast}\setminus D_{F}$ has singleton preimage in the limit $Z$. We thus call any such $p$ the stable point of $K_{F}^{\ast}$. We will casually abuse notation and refer to the preimage in $Z$ or $K_{F^{\prime}}^{\ast}$ of a stable point $p\in K_{F}^{\ast}$ by the same name $p$.
@@
If $K$ is the realization of a graph $\Gamma$ then every edge of $\Gamma$ is an topological edge of $K$.
@@
We will argue that $q\notin C$ so that $U=C\cap\bigl{(}e^{\prime}\setminus\{p,q\}\bigr{)}=C\cap f_{FF^{\prime}}^{-1}(e\setminus\{p,q\})$. For the sake of reaching a contradiction, suppose $C$ contains $q$ in addition to $p$. Let $C^{\prime}$ be the union of the remaining components of $e^{\prime}$. Then $C^{\prime}$ is open in $f_{FF^{\prime}}^{-1}(e\setminus\{p,q\})$, which is open in $K_{F^{\prime}}^{\ast}$, so $C^{\prime}$ is open in $K_{F^{\prime}}^{\ast}$. On the other hand, since $e^{\prime}$ has finitely many components (being homeomorphic to a finite graph), we see that $C^{\prime}$ is closed in $e^{\prime}$, which is closed in $K_{F^{\prime}}^{\ast}$, so $C^{\prime}$ is closed in $K_{F^{\prime}}^{\ast}$. Then $K_{F^{\prime}}^{\ast}$ is disconnected, which contradicts connectedness of $Z$ by Fact 4.5. So we have $U=C\cap f_{FF^{\prime}}^{-1}(e\setminus\{p,q\})$.
@@
Suppose no such $V$ exists. Then, for each $n\in\mathbb{N}$, there exists a point $p_{n}\in X\setminus U$ such that $f(p_{n})\in B_{\frac{1}{n}}(p)$. Since $X$ is compact, it has a convergent subsequence $(p_{n_{k}})_{k}$. The limit $p^{\prime}$ of this subsequence is also contained in the closed set $X\setminus U$. In particular $p^{\prime}\neq\hat{p}$. But, by continuity of $f$, we have $f(p^{\prime})=\lim_{k\to\infty}f(p_{n_{k}})=p$, contradicting $f^{-1}(p)=\{\hat{p}\}$. ∎
@@
The stable points of $K_{F}^{\ast}$ are contained in the interiors of edges so, if an endpoint of $e$ is unstable we can slightly enlarge $e$ to avoid this. Thus, without loss of generality, the endpoints of $e$ are unstable. We will start by proving that $N$ satisfies condition (1) of Definition 5.1. Let $a$ and $b$ be the endpoints of $e$. Let $p$ and $q$ be distinct stable points of $e$. Then $p,q\in e\setminus\{a,b\}$. For points $x,y\in e$ let $(x,y)$ denote the open interval of $e$ bounded by $x$ and $y$. Without loss of generality, the components of $e\setminus\{a,b,p,q\}$ are $(a,p)$, $(p,q)$ and $(q,b)$. By Corollary 5.4, the preimage $f_{F}^{-1}\bigl{(}(p,q)\bigr{)}$ is connected. So, to ensure that $Z\setminus\{p,q\}$ has two components it suffices to prove that $Y=Z\setminus\Bigl{(}f_{F}^{-1}\bigl{(}(p,q)\bigr{)}\cup\{p,q\}\Bigr{)}$ is connected.
For the sake of reaching a contradiction assume that $Y$ is disconnected. Then there exist open subsets $U$ and $V$ of $Z$ such that $Y\subset U\cup V$ and $Y\cap U\neq\emptyset$ and $Y\cap V\neq\emptyset$. Since $Y$ is open in $Z$ we can assume that $U$ and $V$ are subsets of $Y$ so that $U\sqcup V=Y$. By Corollary 5.4, the preimages $f_{F}^{-1}\bigl{(}(p,b)\bigr{)}$ and $f_{F}^{-1}\bigl{(}(q,b)\bigr{)}$ are connected. The latter preimage $f_{F}^{-1}\bigl{(}(q,b)\bigr{)}$ is contained in $Y$ so, without loss of generality, we have $f_{F}^{-1}\bigl{(}(q,b)\bigr{)}\subset V$. The intersection $Y\cap f_{F}^{-1}\bigl{(}(p,b)\bigr{)}$ is equal to $f_{F}^{-1}\bigl{(}(q,b)\bigr{)}\subset V$ so $V^{\prime}=V\cup f_{F}^{-1}\bigl{(}(p,b)\bigr{)}$ is open and disjoint from $U$. But then $U\cup V^{\prime}=Y\cup f_{F}^{-1}\bigl{(}(p,b)\bigr{)}=Z\setminus\{p\}$ which is a contradiction since $Z$ has no cutpoints. Thus $Y$ is connected. This completes the proof of condition (1).
It remains to prove that $N$ satisfies condition (2) of Definition 5.1. Let $p$ be a stable point of $e$. Then $p\in e\setminus\{a,b\}$. Let $(U_{n})_{n\in\mathbb{N}}$ be a descending sequence of open intervals of $e$ that contain $p$ such that $\cap_{n}U_{n}=\{p\}$. By Lemma 5.5 and Corollary 5.4, the sets $\bigl{(}f_{F}^{-1}(U_{n})\bigr{)}_{n}$ form a basis of connected open neighbourhoods of $p\in Z$. Since $U_{n}\setminus\{p\}$ is disconnected and $f_{F}$ is surjective, the preimage $f_{F}^{-1}(U_{n})\setminus\{p\}$ in $Z$ must also be disconnected. Since $U_{n}\setminus\{p\}$ has two components, each of which is an open interval of $e$, the preimage $f_{F}^{-1}(U_{n})\setminus\{p\}$ must have exactly two components. It follows that $Z\setminus\{p\}$ has exactly two ends, which completes the proof of condition (2). ∎
Next results presents some property crucial from the perspective of our application of subnecklaces in the main argument of this subsection (provided at its very end).
@@
The closure of $N$ is the limit set $\partial K$ of a flexible factor $K$ of $G$. Since $H$ and $K$ are stabilizers of vertices of the reduced Bowditch JSJ tree $\mathcal{T}^{r}_{G}$ the intersection $H\cap K$ is contained in an edge stabilizer of $\mathcal{T}^{r}_{G}$. Thus $H\cap K$ is either finite or $2$-ended. Since $H$ and $K$ are both quasiconvex (e.g. by an observation in the paragraph right before Corollary 3.7), we have $\partial H\cap\partial K=\partial(H\cap K)$ as subspaces of $\partial G$ [21, Corollary to Theorem 13]. Thus $\partial H\cap N\subset\partial H\cap\partial K=\partial(H\cap K)$ and so $\partial H\cap N$ has cardinality at most $2$. ∎

### v2

@@ 4.1. Trees of graphs as inverse limits of graphs
induced by the map $\varphi_{[t,u]}\circ\varphi^{-1}_{[s,u]}:\varphi_{[s,u]}(\Sigma_{u})\to\varphi_{[t,u]}(\Sigma_{u})$ (under the canonical identification of each of the two divisors with the corresponding peripheral subspace), and that $b_{t,s}=b_{s,t}^{-1}$. The equivalence relation $\sim$ is then determined by all of the maps $b_{s,t}$ as above related to all white vertices $u$ of $F$. Note that the described quotient space $X_{F}$ may be viewed as obtained by appropriate iterated connected sum of the family $\Gamma_{t}:t\in V^{b}_{F}$ of graphs (see the paragraph right before Lemma 4.8 for a precise description of the concept of connected sum for graphs). Note also that $X_{F}$ is in a natural way homeomomorphic to the space $K^{*}_{F}$, via homeomorphism induced by the homeomorphisms $h_{t}^{-1}:|\Gamma_{t}|\to K^{*}_{t}$, for $t\in V^{b}_{F}$.
In order to make our alternative description of the inverse system ${\mathcal{S}}_{\Theta}$ complete, we need to describe the bonding maps between the spaces $X_{F}$ which, under the above identifications with the spaces $K_{F}^{*}$, correspond to the bonding maps $f_{FF^{\prime}}$. To do this, consider any two finite b-subtrees $F\subset F^{\prime}$ of $T$, and define the map $\pi_{FF^{\prime}}:X_{F^{\prime}}\to X_{F}$ to be a natural map from the larger connected sum to the smaller one, given as follows. For $t\in V^{b}_{F}$, the restriction of $\pi_{FF^{\prime}}$ to $|\Gamma_{t}|^{\#}(J_{t}^{F^{\prime}})$ (viewed as a subset in $X_{F^{\prime}}$) coincides with the blow-down map $\rho^{|\Gamma_{t}|}_{J_{t}^{F^{\prime}}J_{t}^{F}}$ (whose image $|\Gamma_{t}|^{\#}(J_{t}^{F})$ is viewed as a subset in $X_{F}$). For $t\in V^{b}_{F^{\prime}}\setminus V^{b}_{F}$, the subset $|\Gamma_{t}|^{\#}(J_{t}^{F^{\prime}})\subset X_{F^{\prime}}$ is mapped by $\pi_{FF^{\prime}}$ to the point $h_{s}([\varphi_{[u,s]}(\Sigma_{u})])\in|\Gamma_{s}|^{\#}(J_{s}^{F})\subset X_{F}$, where $s\in V^{b}_{F}$ is the vertex of the subtree $F$ that is closest to $t$, and where $u\in V^{w}_{F^{\prime}}$ is the vertex adjacent to $s$ on the shortest path in $T$ from $t$ to $s$.
@@
Expression of a tree of graphs $\lim\Theta$ as the inverse limit of graphs $X_{F}$, as above, turns out to be very useful for establishing various topological properties of trees of graphs. The next corollary records the most obvious such properties (which follow directly from the corresponding well known properties of inverse limits, e.g. the estimate of the topological dimension of the limit by supremum of the topological dimensions of the terms in the inverse system). Further properties, for a subclass of trees of 2-connected graphs, are discussed in the next subsection.
@@
Any tree of graphs is a compact metrizable topological space of topological dimension $\leq 1$.
@@ 4.2. Trees of 2-connected graphs
A graph is 2-connected if it is finite, connected, nonempty, not equal to a single vertex, and contains no cutpoint. We say that a tree system of punctured graphs $\Theta=(T,\{\Gamma_{t}^{\circ}\},\{\Sigma_{u}\},\{\varphi_{e}\})$ in which all the graphs $\Gamma_{t}$ are 2-connected is a tree system of punctured 2-connected graphs; likewise, a tree of 2-connected graphs is the limit of a tree system of punctured 2-connected graphs.
@@
Let ${\mathcal{S}}=(\{K_{\alpha}\}_{\alpha},\{f_{\alpha\alpha^{\prime}}\}_{\alpha\leq\alpha^{\prime}})$ be an inverse system of compact metric spaces with surjective bonding maps. Let $C$ be a closed subspace of some $K_{\alpha}$. Then the preimage $f_{\alpha}^{-1}(C)$ of $C$ in the inverse limit of $\mathcal{S}$ is disconnected if and only if $f_{\alpha\hat{\alpha}}^{-1}(C)$ is disconnected for some $\hat{\alpha}\geq\alpha$.
@@
Let $X$ be a connected, locally connected, cutpoint-free $T_{1}$ space. Let $x,y\in X$ and let $C$ be a component of $X\setminus\{x,y\}$. The closure $\bar{C}$ of $C$ in $X$ is $C\cup\{x,y\}$.
@@
If $x=y$ then we can directly apply Lemma 4.8 so we assume that $x\neq y$. Put $X_{y}=X\setminus\{y\}$ and note that $X_{y}$ is connected (because $X$ is cutpoint-free). Then $C$ is a component of $X_{y}\setminus\{x\}$ so, by Lemma 4.8, the closure of $C$ in $X_{y}$ is $C\cup\{x\}$. Then $\bar{C}$ is either $C\cup\{x\}$ or $C\cup\{x,y\}$. Swapping the roles of $x$ and $y$ we see that $\bar{C}=C\cup\{x,y\}$. ∎
@@
Let $X_{1}$ and $X_{2}$ be realizations of graphs, suppose that $X_{2}$ is 2-connected, and let $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$. Let $\ell\colon P_{x_{1}}\to P_{x_{2}}$ be a bijection between the blow-up divisor $P_{x_{1}}$ of $x_{1}$ in $X_{1}$ and the blow-up divisor $P_{x_{2}}$ of $x_{2}$ in $X_{2}$. Let $A$ be a connected subspace of $X_{1}$. Then the preimage of $A$ under the projection $X_{1}\#_{\ell}X_{2}\to X_{1}$ is connected.
@@
It is clear that $X$ is not empty and not equal to a single point. So we need only to show that $X$ is connected and has no cutpoint. That $X$ is connected follows from Lemma 4.10 by setting $A=X_{1}$.
@@
By Fact 4.5, it suffices to prove that $f_{FF^{\prime}}^{-1}(C)$ is connected for any b-subtree $F^{\prime}\supset F$. We prove this by induction on the size of $F^{\prime}\setminus F$. The base case $F^{\prime}=F$ is immediate. Otherwise, the reduced partial union graph $K_{F^{\prime}}^{\ast}$ is the connected sum of the reduced partial union graph $K_{F^{\prime\prime}}^{\ast}$ and the realization of a 2-connected graph, for some b-subtree $F^{\prime\prime}$ with $F^{\prime}\supsetneq F^{\prime\prime}\supset F$. Then $f_{FF^{\prime}}^{-1}(C)=f_{F^{\prime\prime}F^{\prime}}^{-1}(\hat{C})$ where $\hat{C}=f_{FF^{\prime\prime}}^{-1}(C)$, which is connected by the inductive hypothesis. By Lemma 4.10, the preimage $f_{F^{\prime\prime}F^{\prime}}^{-1}(\hat{C})=f_{FF^{\prime}}^{-1}(C)$ is connected. ∎
@@
Let ${\mathcal{S}}_{\Theta}=(\{K_{F}^{\ast}:F\in{\mathcal{F}}_{T}^{b}\},\{f_{FF^{\prime}}:F\subset F^{\prime}\})$ be the standard inverse system associated to $\Theta$. By Lemma 4.11, any reduced partial union $K_{F}^{\ast}$ is connected. Then, since an inverse limit of connected metric compacta is connected (or by Fact 4.5), $Z$ is connected as well.
@@
whose limit $\lim_{\leftarrow}{\mathcal{C}}$ is obviously a subspace of $Z$. Observe that, since $\mathcal{C}$ is a sequence of circles, with bonding maps which are near-homeomorphisms (i.e. can be approximated by homeomorphisms), by the result of M. Brown [Bro60] the limit $\lim_{\leftarrow}{\mathcal{C}}$ is also a circle, which completes the proof. ∎
5. Necessary conditions for tree of graphs boundary
This section is devoted to the proofs of the implications (1$\Rightarrow$(3) of Theorem A and (1$\Rightarrow$(3) of Theorem B.
5.1. Proof of the impliciation (1$\Rightarrow$(3) of Theorem A
In this subsection we will introduce subnecklaces, prove some results about sufficiently connected trees of graphs and use these results to prove the impliciation (1$\Rightarrow$(3) of Theorem A.
@@ 5.1. Proof of the impliciation (1$\Rightarrow$(3) of Theorem A
Let $X$ be a connected, locally connected, cutpoint-free, compact Hausdorff space. An infinite subspace $N\subset X$ is a subnecklace if it satisfies the following conditions.
@@
    For any $p,q\in N$, the space $X\setminus\{p,q\}$ has two components.
@@
    For any $p\in N$, the space $X\setminus\{p\}$ has two ends (i.e. $p$ is a local cutpoint of degree 2).
@@
We now turn to the analysis of subnecklaces in connected, locally connected and cutpoint-free trees of graphs. We start with terminological preparations, and with some preparatory technical observations. Let $\Theta=(T,\{\Gamma_{t}^{\circ}\},\{\Sigma_{u}\},\{\varphi_{e}\})$ be a tree system of punctured graphs, let ${\mathcal{S}}_{\Theta}=(\{K_{F}^{\ast}:F\in{\mathcal{F}}_{T}^{b}\},\{f_{FF^{\prime}}:F\subset F^{\prime}\})$ be the standard inverse system associated to $\Theta$ and assume that the limit $Z=\lim\Theta=\lim_{\leftarrow}\mathcal{S}_{\Theta}$ is connected and locally connected. As described in Section 4.1, for a finite b-subtree $F$ of $T$, the reduced partial union $K_{F}^{\ast}=K_{F}/\Omega_{F}$ of $\Theta$ is a graph. Let $D_{F}$ be the image of the union of the peripheral subspaces $\Omega_{F}$ in $K_{F}^{\ast}$ and note that $D_{F}$ is dense, countable and contains the vertex set of $K_{F}^{\ast}$. Note that any $p\in K_{F}^{\ast}\setminus D_{F}$ has singleton preimage in the limit $Z$. We thus call any such $p$ the stable point of $K_{F}^{\ast}$. We will casually abuse notation and refer to the preimage in $Z$ or $K_{F^{\prime}}^{\ast}$ of a stable point $p\in K_{F}^{\ast}$ by the same name $p$.
@@
If $K$ is the realization of a graph $\Gamma$ then every edge of $\Gamma$ is a topological edge of $K$.
@@
We will argue that $q\notin C$ so that $U=C\cap\bigl{(}e^{\prime}\setminus\{p,q\}\bigr{)}=C\cap f_{FF^{\prime}}^{-1}(e\setminus\{p,q\})$. For the sake of reaching a contradiction, suppose $C$ contains $q$ in addition to $p$. Let $C^{\prime}$ be the union of the remaining components of $e^{\prime}$. Then $C^{\prime}$ is open in $f_{FF^{\prime}}^{-1}(e\setminus\{p,q\})$, which is open in $K_{F^{\prime}}^{\ast}$, so $C^{\prime}$ is open in $K_{F^{\prime}}^{\ast}$. On the other hand, since $e^{\prime}$ has finitely many components (being homeomorphic to a finite graph), we see that $C^{\prime}$ is closed in $e^{\prime}$, which is closed in $K_{F^{\prime}}^{\ast}$, so $C^{\prime}$ is closed in $K_{F^{\prime}}^{\ast}$. Then $K_{F^{\prime}}^{\ast}$ is disconnected. Since the bonding maps in ${\mathcal{S}}_{\Theta}$ are obviously surjective, this contradicts connectedness of $Z$ by Fact 4.5. So we have $U=C\cap f_{FF^{\prime}}^{-1}(e\setminus\{p,q\})$.
@@
Suppose no such $V$ exists. Then, for each $n\in\mathbb{N}$, there exists a point $p_{n}\in X\setminus U$ such that $f(p_{n})\in B_{\frac{1}{n}}(p)$. Since $X$ is compact, there is a convergent subsequence $(p_{n_{k}})_{k}$. The limit $p^{\prime}$ of this subsequence is also contained in the closed set $X\setminus U$. In particular $p^{\prime}\neq\hat{p}$. But, by continuity of $f$, we have $f(p^{\prime})=\lim_{k\to\infty}f(p_{n_{k}})=p$, contradicting $f^{-1}(p)=\{\hat{p}\}$. ∎
@@
The stable points of $K_{F}^{\ast}$ are contained in the interiors of edges so, if an endpoint of $e$ is stable we can slightly enlarge $e$ to avoid this. Thus, without loss of generality, the endpoints of $e$ are unstable. We will start by proving that $N$ satisfies condition (1) of Definition 5.1. Let $a$ and $b$ be the endpoints of $e$. Let $p$ and $q$ be distinct stable points of $e$. Then $p,q\in e\setminus\{a,b\}$. For points $x,y\in e$ let $(x,y)$ denote the open interval of $e$ bounded by $x$ and $y$. Without loss of generality, the components of $e\setminus\{a,b,p,q\}$ are $(a,p)$, $(p,q)$ and $(q,b)$. By Corollary 5.4, the preimage $f_{F}^{-1}\bigl{(}(p,q)\bigr{)}$ is connected. So, to ensure that $Z\setminus\{p,q\}$ has two components it suffices to prove that $Y=Z\setminus\Bigl{(}f_{F}^{-1}\bigl{(}(p,q)\bigr{)}\cup\{p,q\}\Bigr{)}$ is connected.
For the sake of reaching a contradiction assume that $Y$ is disconnected. Then there exist open subsets $U$ and $V$ of $Z$ such that $Y\subset U\cup V$, $Y\cap U\neq\emptyset$ and $Y\cap V\neq\emptyset$. Since $Y$ is open in $Z$ we can assume that $U$ and $V$ are subsets of $Y$ so that $U\sqcup V=Y$. By Corollary 5.4, the preimages $f_{F}^{-1}\bigl{(}(p,b)\bigr{)}$ and $f_{F}^{-1}\bigl{(}(q,b)\bigr{)}$ are connected. The latter preimage $f_{F}^{-1}\bigl{(}(q,b)\bigr{)}$ is contained in $Y$ so, without loss of generality, we have $f_{F}^{-1}\bigl{(}(q,b)\bigr{)}\subset V$. The intersection $Y\cap f_{F}^{-1}\bigl{(}(p,b)\bigr{)}$ is equal to $f_{F}^{-1}\bigl{(}(q,b)\bigr{)}\subset V$ so $V^{\prime}=V\cup f_{F}^{-1}\bigl{(}(p,b)\bigr{)}$ is open and disjoint from $U$. But then $U\cup V^{\prime}=Y\cup f_{F}^{-1}\bigl{(}(p,b)\bigr{)}=Z\setminus\{p\}$ which is a contradiction since $Z$ has no cutpoints. Thus $Y$ is connected. This completes the proof of condition (1).
It remains to prove that $N$ satisfies condition (2) of Definition 5.1. Let $p$ be a stable point of $e$. Then $p\in e\setminus\{a,b\}$. Let $(U_{n})_{n\in\mathbb{N}}$ be a descending sequence of open intervals of $e$ that contain $p$ such that $\cap_{n}U_{n}=\{p\}$. By Lemma 5.5 and Corollary 5.4, the sets $\bigl{(}f_{F}^{-1}(U_{n})\bigr{)}_{n}$ form a basis of connected open neighbourhoods of $p\in Z$. Since $U_{n}\setminus\{p\}$ is disconnected and $f_{F}$ is surjective, the preimage $f_{F}^{-1}(U_{n}\setminus\{p\})=f_{F}^{-1}(U_{n})\setminus\{p\}$ in $Z$ must also be disconnected. Since $U_{n}\setminus\{p\}$ has two components, each of which is an open interval of $e$, the preimage $f_{F}^{-1}(U_{n})\setminus\{p\}$ must have exactly two components. It follows that $Z\setminus\{p\}$ has exactly two ends, which completes the proof of condition (2). ∎
The next result presents some property crucial from the perspective of our application of subnecklaces in the main argument of this subsection (provided at its very end).
@@
The closure of $N$ is the limit set $\partial K$ of a flexible factor $K$ of $G$. Since $H$ and $K$ are stabilizers of vertices of the reduced Bowditch JSJ tree $\mathcal{T}^{r}_{G}$ the intersection $H\cap K$ is contained in an edge stabilizer of $\mathcal{T}^{r}_{G}$. Thus $H\cap K$ is either finite or $2$-ended. Since $H$ and $K$ are both quasiconvex (e.g. by an observation in the paragraph right before Corollary 3.7), we have $\partial H\cap\partial K=\partial(H\cap K)$ as subspaces of $\partial G$ [Swe01, Corollary to Theorem 13]. Thus $\partial H\cap N\subset\partial H\cap\partial K=\partial(H\cap K)$ and so $\partial H\cap N$ has cardinality at most $2$. ∎